OpenStudy (anonymous):

Find the x intercept.8x+4y=12 MEDAL!!!!!!!

OpenStudy (anonymous):

simple! y=mx+b form can you do that?

OpenStudy (igreen):

Input 0 for y

OpenStudy (anonymous):

or just go straight there^ lol

OpenStudy (igreen):

8x + 4y = 12 8x + 4(0) = 12 8x = 12 x = 1.5

OpenStudy (anonymous):

EWWW! A ZOMBIE!

OpenStudy (anonymous):

thanks both of you i wish i could give out more than one medal

OpenStudy (igreen):

I gave @dsood15 one :)

OpenStudy (anonymous):

and im not a zombie i just like eatting brains ;-;

OpenStudy (anonymous):

What is this thing "Unlock tiles" ??

OpenStudy (anonymous):

*Unlock titles ??

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

??

OpenStudy (anonymous):

How Human Calculator is coming in front of your name?

OpenStudy (anonymous):

Oh go to your profile

OpenStudy (igreen):

@waterineyes Click on your name at the top right corner. You will see a purple button at the top right. It says "unlock medals". Click that then buy titles!

OpenStudy (anonymous):

You have to get a certain amount of medals to get this stuff.

OpenStudy (igreen):

But you should have enough..lol. @waterineyes

OpenStudy (igreen):

@conqueror

OpenStudy (anonymous):

What is this for??

OpenStudy (igreen):

Sorry that was an accident.

OpenStudy (anonymous):

No no, what is the concept of this smartcents??

OpenStudy (igreen):

It's a new thing they had..you get smartcents from medals, fans, testimonials, etc. Smartcents BUY YOU TITLES. There are certain requirements though. -Certain # of Medals earned in one category -Certain # of Smartcents

OpenStudy (anonymous):

@iGreen ur hella smart man can i pm u some math questions?

OpenStudy (igreen):

Lol, um..sure.

OpenStudy (anonymous):

that sounded weird huh

OpenStudy (anonymous):

Okay I got now.., Thanks.>!!!

OpenStudy (igreen):

Np! Surprised you didn't know.

OpenStudy (igreen):

?

OpenStudy (anonymous):

?